Project

General

Profile

Ops/Comm Scheduling meeting

Location : CR22 in R1 @ 15:30 pm on Tuesday 7th

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Actions from last meeting
  • Generate appropriate table to store absorber pressure and material information (Janusz) : Will be addressed on Friday
  • DS readback current requires calibration (DL) : Asked Daresbury when they can schedule it. Possibly next week depending on workload.
  • get a ramp-up plan from Jaroslaw. (Steve, Chris)
  • Josef send FC plan to Colin/DC.
  • Ops Coord/DC to contact Melissa about the Hall probe communication problems : Done
  • look at putting wire gauges onto PRY end plates to properly monitor movement (John/Pierrick)
  • Check that PRY movement monitor measurements are in the archiver (Pierrick) : Stands

Ongoing work

  • SS QD/QP commissioning
  • Electrical work
  • Water system work
  • Conventional magnet test report
  • FC cooldown
  • GVA1 removal
  • Planning for July running
    • Expert startup date : : Monday 27th / Tuesday 28th are expert startup dates. Hall locked on Friday 24th.
    • Beam tuning on Tuesday 28th @ 11am ?
    • Operational issues
      • SSH lockout procedure : tested and working. Need an implementation.
      • Completion and return of spreadsheets informing alarm limits
      • Procedures for shifters
        • Shifter tasks and documentation
        • Emergency procedures and documentation for shifters and experts
        • Procedure whilst control room is unmanned
        • Run Conditions Spreadsheet vs CDB
        • Shift checklist update
    • Expert list
    • Formal freeze of online software
  • AOB
  • DONM : 21/6/2016

Minutes

Present : John, Ray, Ryan, Ken, Pail. Mike, Mark, Victoria, Vicky, John, Josef, Paolo, Andy, Melissa, Pierrick, Paul H

Actions from last meeting
  • Generate appropriate table to store absorber pressure and material information (Janusz) : In progress - Janusz coming on Wed.
  • DS readback current requires calibration (DL) : Asked Daresbury when they can schedule it. Possibly next week depending on workload. : Stands
  • get a ramp-up plan from Jaroslaw. (Steve, Chris) : Stands - but discussion on Tuesday @ 11
  • Josef send FC plan to Colin/DC. : Communicated with Steve Griffiths on steps to follow. John thinks it's under control. Steve G's plan is not quite complete but will be soon.
  • Ops Coord/DC to contact Melissa about the Hall probe communication problems : Done : Ajit looking into it. In progress. There was a problem in one of the Hall probe cables got knocked again - can we shield these? Another issue is that the DHCP connection occasionally drops.
    Action : John to find a shielding solution for the Hall probe connectors.
  • look at putting wire gauges onto PRY end plates to properly monitor movement (John/Pierrick) : Discussed. Know where to put the probes. Will put 2 on the upper east and west wall. In hand. Josef was in communication with Holge and he says that we are well in the linear range. Even 10mm shift would be fine. A 1mm shift is no concern. Whole PRY is slightly floppy so we don't expect movement to relax exactly back into its new position.
  • Check that PRY movement monitor measurements are in the archiver (Pierrick) : They are now in the archiver.

Ongoing work

  • SS QD/QP commissioning : pretty much done. FInished this week - should be ready for Shandor & Alan starting on the 28th.
    • Paul H : what is the plan for powering the magnets in parallel? Mike - need to get the cross-calibration of the currents between the GUIs, DCCTs and power supplies right. Action : Ask Daresbury to check the current calibrations for the FC (John G).
  • Electrical work : Discussed during DC report
  • Water system work : Standstill at the moment. Chiller is running and the water is clean. Want to start putting the TPCs in the separate cryocompressors. This will happen on Thursday.
  • Conventional magnet test report : Summary on the elog. Can probably run the magnets to the required currents. The high momentum settings may require tweaking the cooling. Such tweaks will happen during the day.
  • FC cooldown and testing : Cooldown complete. Testing of FC beginning.
  • GVA1 removal : Done and DSA closed. Whoever is in charge of the G4Beamline should generate new beams without GVA1. Ryan has been informed.
  • Planning for July running
    • Expert startup date : : Monday 27th / Tuesday 28th are expert startup dates. Hall locked on Friday 24th.
    • Beam tuning on Tuesday 28th @ 11am ? Verified with Bryan.
    • Operational issues
      • SSH lockout procedure : tested and working. Need an implementation : needs to go into the startup checklist. document procedure. Paolo will document
        procedure. Henry/I will modify checklist. Action : Paolo will document SSH lockout process.
        Action : Henry/I will think about MOM checklist modification.
      • Completion and return of spreadsheets informing alarm limits : Have all spreadsheets been completed? Need final bits for powering the magnets for
        the solenoids. Decay Solenoid - need to meet with Mark. Focus Coil - will need to go over with John. Pierrick - will need to put an operating alarm
        handler on the computer of each subsystem owner.
        *Action : * Meet with Mark and John Cobb to finalise alarm handler spreasheets for DS and FC.
      • Action : Alarm handler training.
      • John C : something which would be really useful would be to be able to look at the archiver data over the net. Pierrick : RAL network opens
        up certain ports outside RAL networks. We could open up a port from external machines. Josef : I can do this with no problem.
      • Procedures for shifters
        • Shifter tasks and documentation : Andy. For other subsystems we have a comprehensive set of operating instructions. We need a set of documents
          to document how shifters/experts interact with the magnet powering. Ed: Only comment is that there needs to be someone on call for the SC magnets.
          It will be a combination between Josef/Mark. Mike : it will have to be an expert-only system until the documentation is complete. Josef providing
          all documentation.
          *Action : * Keep on top of documentation process (Josef with help from all)
        • Emergency procedures and documentation for shifters and experts : same document
        • Procedure whilst control room is unmanned : kept energised with MCR oversight.
        • Run Conditions Spreadsheet vs CDB : The eternal argument begins. Recommendation : CDB viewer needs to have the cooling channel magnets
          and absorber status. Action : Janusz needs to add cooling channel magnet currents and absorber to the CDB viewer.
          Options for searching through the CDB exist.
          Action Item : Get a search function installed somehow by someone.
        • Shift checklist update : suggestions
    • Expert list
    • Formal freeze of online software : break it down by subsystem. Freeze most systems on Tuesday 28th. Allow one iteration. Then change control.

2016-06-13-Workplan.xlsx (12.7 KB) Hodgson, Paul, 14 June 2016 14:31